OpenStudy (anonymous):

What is the solution to the system of equations y=3x-7 6x-2y=12 A.(12,14) B. infinite number of solutions C. (14,12) D. no solutions

OpenStudy (anonymous):

I think the answer is d

OpenStudy (tkhunny):

Prove it.

OpenStudy (anonymous):

I don't think there are any solutions when I plug the other answers in they don't work

OpenStudy (anonymous):

I'm also not quite sure what to do

OpenStudy (tkhunny):

What did you do? y = 3x - 7 6x - 2y = 12 ==> 3x-y=6==> y = 3x - 6 So, then... y = 3x - 7 y = 3x - 6 In this form it is an eyeball problem. Same slope - parallel. Different intercept. No solution.

OpenStudy (anonymous):

I'm just not sure how to simplify when an equation is in y intercept form

OpenStudy (anonymous):

@tkhunny thanks for the reply

OpenStudy (tkhunny):

Simplify what? One was already in that form. The other was just solved for y.

OpenStudy (anonymous):

oh ok I get it now thanks

OpenStudy (anonymous):

I think all I had to do is simplify the other equation to y intercept form

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Thanks for all the help